OpenStudy (anonymous):

The space shuttle releases a satellite into a circular orbit 520 km above the Earth. How fast must the shuttle be moving (relative to Earth) when the release occurs?

OpenStudy (anonymous):

Centripetal force F= GMm/r^2 which is equal to (mv^2)/r so GMm/r^2 = (mv^2)/r now you can calculate v by putting r =R+520 km R is radius of earth and M is mass of earth G gravitational constant m is mass of space shuttle which will cancel out in your calculation.

OpenStudy (jfraser):

@gyanu don't forget to convert the altitude (520km) into METERS, otherwise it doesn't work properly

OpenStudy (anonymous):

thanks @JFraser but i left it for @ppatel428.

OpenStudy (jfraser):

oh, sorry!